I need with this problem 6y^2+12y+13=2y^2+4
Answer by checkley77(12844)

6y^2+12y+13=2y^2+4
6Y^2-2Y^2+12Y+13-4=0
4Y^2-12Y+9=0
(2Y-3)(2Y-3) OR (2Y-3)^2 ANSWER.